Cyclohexylbenzene has the empirical formula C3H4. Its molar mass is 160.3 g/mol. Its molecular formula is:
Answer #1

Molecular formula = (empirical formula)*n

Also, Molar mass= (Empirical formula mass)*n [Empirical formula mass= [(3*12)+(4*1)]=36+4=40

hence n=Molar mass/Empirical formula mass=160/40=4

therefore the molecular formula is C12H16 (=4*C3H4)